### v1

@@ 5.2 Sorting and Increased Reporting
One finding from Klawitter and Flatt (1998) suggests that LGB workers sort to areas with anti-discrimination laws. If LGB workers were sorting to areas that recently passed anti-discrimination laws, then it could violate the identifying assumption. It would be concerning if high-wage LGB workers migrated from unprotected areas to areas that recently passed anti-discrimination laws since it would lower the average wage in untreated counties and increase the average wage in treated counties.
@@
In this regression, $\beta_{0}$ is the coefficient of interest, and it will give the effect of anti-discrimination laws passing on the number of same-sex partnerships in a given county whether that is from in-migration or increased reporting of same-sex partnerships. We present these results in Table 8.
@@
6 Discussion
6.1 Differences in Response by Sex
The results paint an interesting picture of the effect of anti-discrimination laws on the workforce, in that they have differing effects on lesbian and gay workers. Lesbian workers see a reduction in the labor supply and pay premium while gay workers see a reduction in their labor supply and pay gap. The anti-discrimination laws appear to push the labor market outcome to a greater level of equality across sexual orientation, where lesbian women move toward the labor market outcomes of straight women and gay men move toward the labor market outcomes of straight men.
One potential explanation for this convergence of labor market outcomes is that employers stop differentiating between workers based on sexual orientation after anti-discrimination laws are passed and that results in greater equality. However, much of the labor market differences arise from differences in human capital accumulation and decisions that come before many of the anti-discrimination laws such as occupation/industry choice and desired number of children (Black et al., 2007).
Another explanation for different responses to anti-discrimination laws for same-sex couples lies in the canonical Becker (1981) theory of the family and division of labor, which is typically used to explain differences in LGB labor market outcomes from their straight counterparts (Black et al., 2007). In a male-female household with a “traditional” division of labor, the man specializes in market production, selling his labor for money to buy market goods, and the woman specializes in household production, using her labor to rear children, cook, and take care of the home. However, LGB people are less likely to end up in a traditional Beckerian household. This expectation of future household composition will induce gay men to be less specialized in market production and lesbian women to be more specialized in market production than their straight counterparts. This theory of household specialization is typically used to explain why gay men make less than straight men and lesbian women make more than straight women (Black et al., 2007).
This household model can also be used to think about responses to anti-discrimination laws. One potential explanation for lesbian women becoming more likely to leave the workforce and earn less is that a lesbian household may be more likely to shift to a more traditional Beckerian household where one parent specializes in household production and the other shifts to a more market orientation. If anti-discrimination laws decreased the risk of the primary earner getting fired for their sexual orientation, it could induce the secondary earner to switch to more part-time work or exit the workforce to specialize in household production or focus on child-rearing. Alternatively, both women in a female same-sex partnership could switch to reducing hours to spend more time in household production as a result of increased job security.
Households consisting of lesbian women typically have more children than households consisting of gay men. Given lesbian women have more children than gay men, it would be unsurprising to see lesbian households switch more to a specialization of their household labor than gay men would, which would help explain why anti-discrimination laws have different effects on lesbian women and gay men.
To test the different responses by men and women in same-sex partnerships to anti-discrimination laws and the implications of Becker’s household model, we implement an alternative difference-in-difference model. Instead of comparing those in same-sex partnerships to those in different-sex partnerships, we compare men and women in same-sex partnerships to examine their differential responses at the household level. Do women in same-sex partnerships change their household division of labor relative to men in same-sex partnerships following anti-discrimination laws? To answer this question, we collapse the data to the household level and examine households instead of individuals.
Specifically, we estimate this model:
$y_{jct}=\alpha_{0}Law_{ct}+\alpha_{1}FemSSP_{j}+\alpha_{2}Law_{ct}*FemSSP_{j}+\mu_{c}+\delta_{t}+\epsilon_{jct}$
The notation is the same as before with two main differences. Our interaction term uses an indicator for if household $j$ is a female same-sex partnership, denoted by $FemSSP_{j}$. $y_{jct}$ is a variable for outcomes at the household level. $\alpha_{2}$ is our coefficient of interest and will show how lesbian households differ in their response to anti-discrimination laws compared to gay households. The first outcome examined is an indicator for if the household is a one-earner family. We also examine the difference in absolute value between the two partners in terms of weekly hours worked. These labor supply measures inform how intra-household labor is divided between partners along the extensive and intensive margin of labor supply. Finally, we examine how the anti-discrimination laws may affect child rearing by examining if the households have any children and the number of children they have.
We present the results in Table 9. Anti-discrimination laws do not affect the likelihood of same-sex partnerships becoming a one-earner household, but there is an effect on the intensive margin. There is a significant effect of anti-discrimination laws in the intra-household difference in hours worked for gay and lesbian households. An increase in the difference of hours worked within the household could suggest a greater specialization in market production for one person and a greater specialization in household production and working fewer hours for the other person in the partnership. Lesbian households do see a significant increase in the likelihood of having any child and the number of children they have relative to gay households.
This evidence tied with Becker’s theoretical work suggests that men and women in same-sex partnerships are responding differently to these anti-discrimination laws. Lesbian households begin to have more children than gay households. It may appear in the main results that anti-discrimination laws hurt lesbian women because their labor supply and pay premium over straight women shrinks. However, their response in the labor market could be driven by a greater desire for children and subsequent changes in household labor allocation due to an increase in job security. The corresponding changes in the labor market could represent an increase in welfare for lesbian households if they value their new children more than their lost hours worked and subsequently lost wages. Consistent with the theory and empirical findings, lesbian households could become more secure with one woman being the primary earner following anti-discrimination laws.
6.2 Mechanisms
The results of this research indicate that sexual orientation anti-discrimination laws could significantly impact the labor market gaps for gay/bisexual men as well as the labor market premiums for lesbian/bisexual women. It is possible that these laws are binding and eliminate significant discrimination in the labor market. However, it is relatively easy for an employer to create a reason to fire an employee rather than firing someone specifically for being gay. It is also difficult to accurately determine someone’s sexual orientation by looking at them, unlike race and sex. Although it may be easier to determine someone’s sexual orientation conditional on them being in a same-sex relationship. Sexual orientation anti-discrimination laws may not be binding in a meaningful way that eliminates discrimination.
One mechanism that the laws could be affecting labor market outcomes is through increased positive sentiment toward LGB workers. We show in Figure 5 that following state anti-discrimination laws there is a significant and persistent increase in the percentage of people in that state that support same-sex marriage. Our regression results already condition on the change in the county-level support for same-sex marriage via fixed effects. In Appendix A, we show that local laws matter significantly, so it seems plausible that local law changes could significantly change local sentiment and impact the labor supply and pay of LGB workers. Changing public sentiment in conjunction with greater protection for LGB workers seems more plausible as the mechanism than simply greater protection in the workplace.
6.3 Threats to External Validity
One limitation of this research is that we are unable to view all LGB workers. The analysis is specifically conducted on individuals in same-sex partnerships and different-sex partnerships. The data does not allow for the identification of single LGB workers, whose outcomes may be significantly different than LGB workers in partnerships. Another pitfall to using partnerships to infer sexual orientation is that one could incorrectly infer someone’s sexual orientation from a partnership. Bisexual people are a larger group than lesbian or gay people, and this specification could erroneously assign bisexual people to be “straight” through being in a different-sex partnership (Gates, 2011). If bisexual workers appear more similar to their lesbian and gay counterparts then the results of quantifying the labor supply and pay gap/premium would likely be biased toward zero and the change in those labor market gaps/premiums would also be biased toward zero.
7 Conclusion
This analysis is the first quasi-experimental research examining how both local and state anti-discrimination laws on sexual orientation affect the labor supply and pay gap/premium between LGB and straight workers. We construct a novel panel dataset on local anti-discrimination laws to capture the granular nature of anti-discrimination laws. We find that anti-discrimination laws appear to decrease sexual orientation inequality in the labor market. Specifically, we find that anti-discrimination laws significantly reduce the gay labor force participation gap by 2.1 p.p., the employment gap by 1.5 p.p., and increase hourly wages by 6.2%. Anti-discrimination laws also reduce the lesbian labor force participation premium by 1.7 p.p., and the employment premium by 2.3 p.p. We also show that one potential mechanism that anti-discrimination laws work through is increasing positive sentiments toward LGB Americans as measured by the support of same-sex marriage. Finally, we explain the differential response to these laws between gay and lesbian households through Becker’s household specialization model and support the explanation with empirical evidence.
Before 2020, it was legal in some states to fire someone solely based on their sexual orientation, and states have been less inclined to pass anti-discrimination laws recently. Arkansas and Tennessee even passed anti-anti-discrimination laws preventing cities from enacting anti-discrimination laws. From 2009 to 2019, more states have barred cities from protecting their residents than the number of states extending protection. Our research gives a comprehensive look at a policy that Congress is considering in “The Equality Act” to extend federal protection to sexual orientation and gender identity. Our research gives the most comprehensive look at policy and can meaningfully inform the policy discussion around anti-discrimination laws. Overall, this research broadly informs the policy decision of extending employment protections to the LGB community by the Supreme Court of the United States.
Appendix A TWFE Estimates
In the original version of our paper, we employed a traditional difference-in-differences (DiD) setup, incorporating a range of controls to assess the impact of anti-discrimination legislation on labor market outcomes for LGB workers. This conventional approach served as the foundational analysis for our initial findings. To ensure transparency and thoroughness in our research, we are including these initial estimates in this section. Furthermore, in response to recent advancements in DiD methodologies, we use Borusyak et al. [2021] to estimate the impact of anti-discrimination laws. Our analysis reveals that the application of these modern DiD techniques does not substantially alter the point estimates we initially obtained. This consistency across methodological frameworks underscores the robustness of our findings, reinforcing the reliability of our conclusions regarding the effects of anti-discrimination laws on LGB labor market outcomes.
A.1 Identification Strategy & Estimation
The differential roll-out of anti-discrimination laws by state and locality over time lends itself to a difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay between two groups, LGB workers and heterosexual workers. This strategy uses the variation presented in Figure 1, which shows how the laws changed over time by state and county.
Formally, we estimate this equation on labor supply:
$LS_{ict}=\alpha_{0}Law_{ct}+\alpha_{1}SSP_{i}+\alpha_{2}Law_{ct}\cdot SSP_{i}+\alpha_{3}X_{i}+\alpha_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (4)
Where $LS_{ict}$ is the labor force participation or employment status for person $i$ in county $c$ in year $t$. $Law_{ct}$ is an indicator for if county $c$ has an employment anti-discrimination law on sexual orientation or is covered by a state-wide anti-discrimination law in year $t$, and $SSP_{it}$ is an indicator for if person $i$ is in a same-sex partnership. $X_{i}$ is a vector of person-specific covariates including race, children, along with age and education level - each represented up to a quadratic term to capture potential nonlinear effects. $\gamma_{st}$ is a vector for state $s$ covariates on the legality of same-sex marriage in year $t$ as well as polling for the support of same-sex marriage. It includes the percentage of people that “strongly support,” “support,” “oppose,” and “strongly oppose” same-sex marriage in a given state in a given year. Finally, $\mu_{c}$ gives county fixed effects, $\delta_{t}$ gives year fixed effects, and $\epsilon_{itcs}$ is the error term [^15] Since the ACS only identifies metro counties, the county fixed effect is close to a pseudo-county fixed effect. The rural counties are all compared together with a true county fixed effect for metro area counties.. Our sample is limited to people aged 25 to 65, and we estimate this model separately by sex using the ACS person weights. We cluster the standard errors at the county level, as that is the level of the treatment in this specification. [^16] Treatment status is computed using the earliest year of ENDAs at the county or state level, with individuals remaining classified as “treated” thereafter throughout the dataset. Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al. [2020], there’s a potential for bias in treatment assignment. This extended timeframe might not precisely align with the exact implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero. The coefficients of interest are $\alpha_{2}$, which will give the effect of the law on the labor supply gap/premium, and $\alpha_{1}$, which gives the labor supply gap/premium in the absence of anti-discrimination law. Our specification assumes that the effect of the laws is the same across states and localities, and we will be estimating an average effect. We estimate a similar equation for wages:
$\ln(Y_{ict}+1)=\beta_{0}\text{Law}_{ct}+\beta_{1}\text{SSP}_{i}+\beta_{2}\text{Law}_{ct}\cdot\text{SSP}_{i}+\beta_{3}X_{i}+\beta_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (5)
Where $Y_{ict}$ is the real value, in 1999 dollars, of person $i$’s wages in county $c$ year $t$. The rest of the notation is the same as above. The coefficients of interest are $\beta_{2}$, which will give the effect of the law on the pay gap/premium, and $\beta_{1}$, which gives the pay gap/premium. [^17] The $\ln(Y_{ict}+1)$ specification is adopted to include both employed and unemployed individuals in the analysis, recognizing employment as endogenous to anti-discrimination laws. This approach ensures a comprehensive evaluation of the law’s impact on the entire income spectrum, capturing nuances missed by focusing solely on employed workers. We report the coefficients estimated via Borusyak et al. [2021] in the main text, and standard TWFE in the Appendix.
A.2 TWFE Results
In Table A1, we present the regression on the extensive margin of labor supply with Panel A giving the effect of anti-discrimination laws on gay men’s labor supply and Panel B giving the effect for lesbian women. We present the full specification in columns (1) and (2) and include a specification without accounting for local anti-discrimination laws in columns (3) and (4) to show how failing to account for local laws would lead to incorrect inference.
Consistent with the previous literature, we find that lesbian women have a labor supply premium and gay men have a labor supply penalty on the extensive margin. Lesbian women are 7.3% more likely to be employed and gay men are 7.9% less likely to be employed than their straight peers. Anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for gay men. The labor force participation rate gap is reduced by 1.3 p.p. (18%), and the employment gap is reduced by 1.4 p.p. (18%). Conversely, the labor force participation and employment premium that lesbian women have over their straight counterparts is reduced by 1.3 p.p. (16%) and 1 p.p. (14%), respectively. As seen in columns (3) and (4) of Panel A, estimating the effect of anti-discrimination laws only using state laws will lead one to erroneously conclude that these laws have no significant effect on the gay labor force and employment gap. However, the effect on the extensive margin of labor supply for lesbian women looks similar when estimating the effect of just state anti-discrimination laws.
In Table A2, we present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked. Similar to Table 4, we include the results when only estimating the effects of the state laws in columns (3) and (4) to show the contribution of including local laws in estimating the effect of anti-discrimination laws.
We replicate past findings where gay men supply significantly less labor and lesbian women supply significantly more labor on the intensive margin. As with Table A1, we find contrasting effects of anti-discrimination laws on male and female outcomes. We find that the intensive labor supply gap for gay men as measured by weekly hours worked and weeks worked is significantly reduced, and the intensive labor supply premium for lesbian women is significantly reduced as well. The effect of anti-discrimination laws on weekly hours worked becomes insignificant for gay men and lesbian women when only estimating off of state-wide anti-discrimination laws.
In Table A3, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings. These variables received a log-transformation, to handle zero values so the changes can be interpreted in percentage terms. We run the regressions separately by sex and present the regressions showing how the sexual orientation anti-discrimination laws impact both employed and unemployed people in columns (1) and (2) and limit the regressions to only those who are employed in columns (3) and (4). Our preferred specification is to look at both employed and unemployed people in the sample as employment is endogenous. Conditioning on employed workers would create an endogeneity problem since labor force participation and employment are significantly impacted by anti-discrimination laws. Given Table A1, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition. We present the regressions limited to employed workers to replicate past work on the wage gap/premium where it is common to only examine employed workers. We include results by only using state anti-discrimination laws, excluding local laws. We present those results in columns (5) - (8).
Gay men see a significant decline in the hourly wage gap by 2.8 p.p. (11%) at the 90% significance level and see an insignificant decline in the annual earnings gap by 5.8 p.p. On the other hand, lesbian women see a significant change in their annual earnings premium by 13 p.p. (16%) and an insignificant reduction in their hourly wage premium. Columns (3) and (4) replicate the past literature where employed gay men make about 8% to 11% less than their straight counterparts and employed lesbian women make about 5% to 15% more than their straight counterparts.
The pay gap/premium is reduced quite significantly once the sample is limited to employed workers, which is unsurprising. As shown in Table A1 and Table A2, there are significant differences in labor supply for LGB and straight workers. Gay men and straight women are far more likely to be unemployed, out of the labor force, and work fewer hours and weeks, so including their zero wages in the regression will increase the gap/premium.
There is a significant decrease in the hourly wage gap now and an insignificant decrease in the annual wage gap. There is also a significant decrease in the annual earnings premium for lesbian women but no change in their hourly wage premium. This result is consistent with the labor supply results in Table A1 and A2 where their intensive and extensive margin of labor supply premium are declining, so the annual earnings decline, but hourly wage does not.
A.3 Occupation Fixed Effects
We show that even when controlling for occupation there is a significant change in relative labor supply for men and women in same-sex partnerships and a significant change in relative pay for women in same-sex partnerships, indicating there is a significant effect of these laws even within occupation. We present the extensive labor supply and pay regressions with occupation fixed effects in Table A4 and A5.
In these tables, there is a smaller pay gap for gay men and a smaller pay premium for lesbian women when controlling for occupation or industry. However, qualitatively the results seem to be similar to the main results in Table A1 and Table A3 . Gay men see a significant reduction in the labor supply gap, and lesbian women see a significant reduction in their labor supply and annual wage premium. There is no longer a significant effect on the wage gap for gay men. These findings suggest that anti-discrimination laws have a significant effect even within occupations. The effect does not come from workers shifting into occupations with less discrimination, but even when conditional on occupation choice, there is a significant reduction in differences in labor supply across sexual orientations for men and women.
Appendix Tables
@@ A.3 Occupation Fixed Effects
Table A1: Effect of Anti-Discrimination Laws: Extensive Margin of Labor Supply
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium. Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A2: Effect of Anti-Discrimination Laws: Intensive Margin of Labor Supply
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium. Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A3: Effect of Anti-Discrimination Laws: Wages and Earnings
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. Columns (3) and (4) are limited to those who are employed. All of the are in terms of 1999 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent. The first row of coefficients show the effect of anti-discrimination laws on the pay gap or premium, and the second row of coefficients give the pay gap or premium. Columns (5) - (8) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients show the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients give the labor supply gap or premium. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A5: Anti-Discrimination Laws on Pay with Occupation FEs
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects. All of the are in terms of 1999 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent. The first row of coefficients show the effect of anti-discrimination on the pay gap or premium, and the second row of coefficients give the pay gap or premium. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A6: Sorting and Reporting
(1) (2) (3)
All SSP Male SSP Female SSP
Laws 0.000170 0.000633 -0.000304
(0.000489) (0.000572) (0.000573)
Observations 12,872,572 6,295,028 6,577,544
R-squared 0.068 0.072 0.065
  • Notes: Data comes from the 2005-2016 yearly ACS seeing how the number of same-sex partnerships change in a county after the passage of an anti-discrimination law. The first column looks at both men and women with the next two columns separating the sexes. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A7: Differential Responses between Male and Female Same-Sex Partnerships
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing women in same-sex partnerships to men in same-sex partnerships. The first column examines if the household has only one-earner. The second column tests the difference in absolute terms of hours worked between the two partners. The third and fourth columns examines how households differ with having children. The first row of coefficients show the differential effect of anti-discrimination between lesbian and gay households. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.

### v2

@@ A.3 Occupation Fixed Effects
One finding from Klawitter and Flatt [1998] suggests that LGB workers sort to areas with anti-discrimination laws. If LGB workers were sorting to areas that recently passed anti-discrimination laws, then it could violate the identifying assumption. It would be concerning if high-wage LGB workers migrated from unprotected areas to areas that recently passed anti-discrimination laws since it would lower the average wage in untreated counties and increase the average wage in treated counties.
@@
In this regression, $\beta_{0}$ is the coefficient of interest, and it will give the effect of anti-discrimination laws passing on the number of same-sex partnerships in a given county whether that is from in-migration or increased reporting of same-sex partnerships. We present these results in Table O6.
@@
Online Appendix Tables
@@
Table O1: Effect of Anti-Discrimination Laws: Extensive Margin of Labor Supply
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium. Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O2: Effect of Anti-Discrimination Laws: Intensive Margin of Labor Supply
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium. Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O3: Effect of Anti-Discrimination Laws: Wages and Earnings
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. Columns (3) and (4) are limited to those who are employed. All of the are in terms of 2019 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent. The first row of coefficients show the effect of anti-discrimination laws on the pay gap or premium, and the second row of coefficients give the pay gap or premium. Columns (5) - (8) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O4: Anti-Discrimination Laws on Labor Supply with Occupation FEs
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients show the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients give the labor supply gap or premium. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O5: Anti-Discrimination Laws on Pay with Occupation FEs
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects. All of the are in terms of 2019 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent. The first row of coefficients show the effect of anti-discrimination on the pay gap or premium, and the second row of coefficients give the pay gap or premium. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O6: Sorting and Reporting
(1) (2) (3) (4) (5)
All SSP Male SSP Female SSP Age Years of Education
ENDA -0.0002 0.0004 -0.0008 0.0257 -0.0212
(0.0005) (0.0005) (0.0005) (0.0648) (0.0182)
Observations 16,071,420 7,850,821 8,220,599 16,071,420 16,071,420
Mean of DV 0.0122 0.0123 0.0122 45.5220 13.8539
$R^{2}$ 0.0043 0.0081 0.0024 0.0101 0.0467
  • Notes: Data comes from the 2005-2019 yearly ACS seeing how the number of same-sex partnerships change in a county after the passage of an anti-discrimination law. The first column looks at both men and women with the next two columns separating the sexes. Column (4) investigates age and (5) looks at years of education. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table O7: Differential Responses between Male and Female Same-Sex Partnerships
@@
  • Notes: Data comes from the 2005-2016 yearly ACS comparing women in same-sex partnerships to men in same-sex partnerships. The first column examines if the household has only one-earner. The second column tests the difference in absolute terms of hours worked between the two partners. The third and fourth columns examines how households differ with having children. The first row of coefficients show the differential effect of anti-discrimination between lesbian and gay households. Standard errors are clustered at the county level. $p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.